# Adaptive Designs in Fast-Track Registration Processes

### v1

@@ 4.2 Adaptive Study Designs for Fast-Track Processes
We will start with a brief recap of confirmative adaptive designs. Confirmatory (two-stage) adaptive designs are organized into two sequential stages of data collection, with an interim analysis at the end of the first stage (see, for instance, [10] or [18]). The interim data can be used to decide whether to terminate or continue the trial. The study can be terminated due to futility (with an acceptance of $H_{0}$) or due to success, i.e. early rejection of a hypothesis. So far, adaptive study designs play an important role in clinical drug development (see, for example, [19] or [20]). They allow the user to learn from the data of an ongoing trial by performing design modifications at interim when the trial is not terminated. Thus, adaptive designs can be considered as extensions of group sequential designs (see, for instance, [10]). Possible design modifications at interim include sample size re-estimation, selection of a primary endpoint, or dropping of treatment arms in combined phase II/III trials while controlling the type I error rate. Moreover, the data from both stages can be used for the final confirmatory hypothesis test. That means, in contrast to the common situation of an external pilot study that prepares an independent confirmatory study, adaptive designs permit the use of the pilot data for the final confirmatory hypothesis test. As we will see in this section, in the context of fast-track procedures, the use of adaptive designs can increase efficiency substantially in comparison to the common design with two separate stages.
@@
There are various strategies for constructing adaptive tests, for instance, the combination test approach from [21] and [22] and the conditional error function approach from [23]. There is also a direct relationship between them; see, for instance, Section 6.3.2 in [10]. In this paper, we work with the conditional error function approach. The principle is as follows: Significance levels with $0\leq\alpha_{1}<\alpha_{0}\leq 1$ for early rejection and acceptance, respectively, may be predefined. Furthermore, a non-decreasing function $A:\mathbb{R}\rightarrow[0,1]$, referred to as the conditional error function must be specified in advance. Let $Z_{k}=\hat{\vartheta}_{k}\sqrt{I_{k}}\sim\mathcal{N}(\vartheta\sqrt{I_{k}},1)$ denote the z-statistic calculated from the stage $k$ data for testing a hypothesis $H_{0}$. If after the first stage $Z_{1}\geq\Phi^{-1}(1-\alpha_{1})$, the trial is stopped and $H_{0}$ is rejected. If $Z_{1}<\Phi^{-1}(1-\alpha_{0})$, the trial is stopped due to futility with an acceptance of $H_{0}$. In the case $\Phi^{-1}(1-\alpha_{0})\leq Z_{1}<\Phi^{-1}(1-\alpha_{1})$ the trial is continued to stage 2 and then $H_{0}$ is rejected iff $Z_{2}\geq\Phi^{-1}(1-A(Z_{1}))$. With our assumptions under $H_{0}$, the distribution of $Z_{1}$ and the conditional distribution of $Z_{2}$ given $Z_{1}$ are stochastically smaller than or equal to the standard normal distribution. Thus, if the trial is continued to the second stage, the conditional rejection probability is less than or equal to $A(Z_{1})$. If $Z_{1}\geq\Phi^{-1}(1-\alpha_{1})$ is observed after the first stage, the conditional rejection probability $A(Z_{1})$ can be set to $1$ and in the case of a stopping the trial due to a binding futility stopping rule to 00. The Type I error rate for $H_{0}$ is under control if this conditional error function fulfils
$\int_{-\infty}^{+\infty}\,A(z_{1})\phi(z_{1})\mathrm{d}z_{1}=\alpha_{1}+\int_{\Phi^{-1}(1-\alpha_{0})}^{\Phi^{-1}(1-\alpha_{1})}A(z_{1})\phi(z_{1})\mathrm{d}z_{1}\leq\alpha\leavevmode\nobreak\ .$ (9)
@@
For a flexibility with regard to the futility stopping, stopping with $Z_{1}<\Phi^{-1}(1-\alpha_{0})$ can be ignored by not stetting $A$ to 0 in this case, or equivalently, replacing the lower stopping boundary with $\infty$- ∞ in (9). The study can still be stopped with acceptance of $H_{0}$ without inflating the type I error rate. A futility level $\alpha_{0}<1$ that is ignored in (9) is often called non-binding futility level. The flexibility with a non-binding futility boundary comes with the price of a reduced power (if the futiltity stop is executed). In this paper, we consider scenarios with binding (Section 4) and non-binding futility (Section 5) stopping.
@@
The Type I error rate is still under control if the design for the first stage is determined or modified after the first stage based on the interim data, given that the conditional distribution of the resulting second stage z-statistic $Z_{2}$ given $Z_{1}$ is stochastically smaller than or equal to the standard normal distribution under $H_{0}$. This follows from what is sometimes denoted as conditional invariance principle; see e.g. [24].
@@
One example of a design modification is adjusting the second stage sample size or more general, the second stage information. A typical goal of this is to achieve a certain (sufficiently) high conditional probability of rejecting $H_{0}$ when the trial is continued to the second stage, i.e. a certain conditional power. A high conditional power corresponds to a high chance for being successful at the second stage. Thus, in the context of a conditional DiGA registration, a high conditional power at end of the first stage should be a convincing argument for a successful subsequent stage in proving that the DiGA has a positive healthcare effect. It thereby meets the corresponding requirement mentioned in the DiGA guideline [8, pp. 109-110].
@@
The conditional power is the conditional probability to reject $H_{0}$ in the second stage conditional on the first stage data. For the true effect $\vartheta$, the conditional power is given by
$\mathbb{P}_{\vartheta}\left(Z_{2}\geq\Phi^{-1}(1-A(z_{1}))\right)=1-\Phi\left(\Phi^{-1}(1-A(z_{1}))-\vartheta\sqrt{I_{2}}\right)\leavevmode\nobreak\ .$
@@
Since the true effect is unknown, the sample size or information of the second stage can be calculated such that a conditional power of $1-\beta$ is achieved at the observed estimate $\hat{\vartheta}_{1}$ or noncentrality parameter $z_{1}$ of the first stage. There are other strategies for adjusting the information used; see, for example, [10, Section 7.4].
@@
We will now describe how two-stage adaptive designs can be used in the two-stage fast-track procedures. For the application for conditional registration, the data from the first stage is used, and for the application for permanent registration, the data from the first and second stage is used. We consider the case where there is no possibility of an early rejection after the first stage, which means we set $\alpha_{1}:=0$. We demand that the conditional error function $A$ fulfills $A(z_{1})\leq 0.5$ for all $z_{1}\in\mathbb{R}$. This ensures that $H_{0}$ can only be rejected after the second stage when $Z_{2}\geq 0$, i.e. a non-negative effect $\hat{\vartheta}_{2}\geq 0$ is observed. After the second stage, the hypothesis $H_{0}$ is rejected iff $Z_{2}\geq\Phi^{-1}(1-A(Z_{1}))$. We will use two different kinds of conditional error functions. The first one is a slight modification from the conditional error function of the inverse normal method from [25] with equal weights. The second considered conditional error function is a slight modification of the conditional error function of Fisher’s product test from [26] (proposed in [21] and [22] for use in adaptive designs): We define for $z_{0}\in\mathbb{R}\cup\{-\infty\}$
@@
$A_{F,z_{0}}(z_{1}):=\begin{cases}\frac{c_{F}(z_{0})}{1-\Phi(z_{1})}\wedge 0.5,&\text{if $z_{1}\geq z_{0}$}\\ 0,&\text{otherwise}\end{cases}\leavevmode\nobreak\ .$ (11)
The constants $c_{I}(z_{0})$ and $c_{F}(z_{0})$ are chosen such that they are nonnegative and fulfill (with $A:=A_{I,z_{0}}$ or $A:=A_{I,z_{0}}$) condition (9). The chosen $z_{0}$ can be interpreted as a binding futility bound on the z-scala with corresponding early stopping level $\alpha_{0}=1-\Phi(z_{0})$. If there is no possibility for permanent registration after a non-successful conditional registration, then the binding futility boundary $z_{0}=z_{f}$ is a logical choice. This implies that the considered conditional error functions depend on $t_{\xi}(I_{1})$, $\xi$ (and the fixed $\eta_{f}$) because $z_{f}$ depend on these parameters.
@@
$I_{2}^{A}(z_{1}):=I_{2}^{A}(z_{1},I_{2,\min},I_{1}):=\max\left\{I_{2,\min},{I_{1}\left\{\Phi^{-1}(1-\beta)+\Phi^{-1}\big{(}1-A(z_{1})\big{)}\right\}^{2}}/{z_{1}^{2}}\right\},\quad z_{1}\geq z_{f}$ (12)
@@
where the minimum information $I_{2,\min}$ remains to be defined. Because $H_{0}$ is tested at the second stage by a z-test at level $A(z_{1})$ we obtain the overall power
@@
$\mathcal{S}_{A}(I_{2,\min},I_{1})$ $:=\mathbb{P}_{\delta}\left(Z_{1}\geq z_{f},\text{reject\leavevmode\nobreak\ }H_{0}\right)$

### v3

@@ 4.2 Adaptive Study Designs for Fast-Track Processes
We will start with a brief recap of confirmative adaptive designs. Confirmatory (two-stage) adaptive designs are organized into two sequential stages of data collection, with an interim analysis at the end of the first stage (see, for instance, the works of Wassmer and Brannath [15] and Bretz et al [24]). The interim data can be used to decide whether to terminate or continue the trial. The study can be terminated due to futility (with an acceptance of $H_{0}$) or due to success, i.e. early rejection of a hypothesis. So far, adaptive study designs play an important role in clinical drug development (see, for example, the works of Pallmann et al [25] and Cerqueira et al [26]). They allow the user to learn from the data of an ongoing trial by performing design modifications at interim when the trial is not terminated. Thus, adaptive designs can be considered as extensions of group sequential designs [15]. Possible design modifications at interim include sample size re-estimation, selection of a primary endpoint, or dropping of treatment arms in combined phase II/III trials while controlling the type I error rate. Moreover, the data from both stages can be used for the final confirmatory hypothesis test. That means, in contrast to the common situation of an external pilot study that prepares an independent confirmatory study, adaptive designs permit the use of the pilot data for the final confirmatory hypothesis test. As we will see in this section, in the context of fast-track procedures, the use of adaptive designs can increase efficiency substantially.
@@
There are various strategies for constructing adaptive tests, for instance, the combination test approach introduced by Bauer [27] and Bauer and Köhne [28] and the conditional error function approach introduced by Proschan and Hunsberger [29]. There is also a direct relationship between them [15]. In this paper, we work with the conditional error function approach. The principle is as follows: Significance levels with $0\leq\alpha_{1}<\alpha_{0}\leq 1$ for early rejection and acceptance, respectively, may be predefined. Furthermore, a non-decreasing function $A:\mathbb{R}\rightarrow[0,1]$, referred to as the conditional error function must be specified in advance. Let $Z_{k}=\hat{\vartheta}_{k}\sqrt{I_{k}}\sim\mathcal{N}(\vartheta\sqrt{I_{k}},1)$ denote the z-statistic calculated from the stage $k$ data for testing a hypothesis $H_{0}$. If after the first stage $Z_{1}\geq\Phi^{-1}(1-\alpha_{1})$, the trial is stopped and $H_{0}$ is rejected. If $Z_{1}<\Phi^{-1}(1-\alpha_{0})$, the trial is stopped due to futility with an acceptance of $H_{0}$. In the case $\Phi^{-1}(1-\alpha_{0})\leq Z_{1}<\Phi^{-1}(1-\alpha_{1})$ the trial is continued to stage 2 and then $H_{0}$ is rejected iff $Z_{2}\geq\Phi^{-1}(1-A(Z_{1}))$. With our assumptions under $H_{0}$, the distribution of $Z_{1}$ and the conditional distribution of $Z_{2}$ given $Z_{1}$ are stochastically smaller than or equal to the standard normal distribution. Thus, if the trial is continued to the second stage, the conditional rejection probability is less than or equal to $A(Z_{1})$. If $Z_{1}\geq\Phi^{-1}(1-\alpha_{1})$ is observed after the first stage, the conditional rejection probability $A(Z_{1})$ can be set to $1$ and in the case of stopping the trial due to a binding futility stopping rule to 0. The Type I error rate for $H_{0}$ is under control if this conditional error function fulfills
$\int_{-\infty}^{+\infty}\,A(z_{1})\phi(z_{1})\mathrm{d}z_{1}=\alpha_{1}+\int_{\Phi^{-1}(1-\alpha_{0})}^{\Phi^{-1}(1-\alpha_{1})}A(z_{1})\phi(z_{1})\mathrm{d}z_{1}\leq\alpha\penalty 10000\ .$ (10)
@@
For flexible decision making with regard to the futility stopping, stopping with $Z_{1}<\Phi^{-1}(1-\alpha_{0})$ can be ignored by not setting $A$ to 0 in this case, or equivalently, replacing the lower stopping boundary with -$\infty$ in (10). The study can still be stopped with acceptance of $H_{0}$. A futility level $\alpha_{0}<1$ that is ignored in (10) is often called non-binding futility level, and the resulting flexibility comes with the price of reduced power (if the futility stop is executed). In this paper, we consider scenarios with binding (Section 4) and non-binding (Section 5) futility stopping.
@@
The Type I error rate is still under control if the design for the first stage is determined or modified after the first stage based on the interim data, given that the conditional distribution of the resulting second stage z-statistic $Z_{2}$ given $Z_{1}$ is stochastically smaller than or equal to the standard normal distribution under $H_{0}$. This follows from what is sometimes denoted as conditional invariance principle; see e.g. the work of Brannath et al [30].
@@
One example of a design modification is adjusting the second stage sample size, or more generally, the second stage information. A typical goal of this is to achieve a sufficiently high conditional probability, i.e. a certain conditional power, of rejecting $H_{0}$ when the trial is continued to the second stage. Thus, in the context of a conditional DiGA registration, a high conditional power at the end of the first stage should be a convincing argument for a successful subsequent stage in proving that the DiGA has a positive healthcare effect. It thereby meets the corresponding requirement mentioned on pages 109-110 in the DiGA guideline [12].
@@
The conditional power is the conditional probability of rejecting $H_{0}$ in the second stage, conditional on the first stage data. For the true effect $\vartheta$, the conditional power is given by
$\mathbb{P}_{\vartheta}\left(Z_{2}\geq\Phi^{-1}(1-A(z_{1}))\right)=1-\Phi\left(\Phi^{-1}(1-A(z_{1}))-\vartheta\sqrt{I_{2}}\right)\penalty 10000\ .$
@@
Since the true effect is unknown, the sample size or information of the second stage can be calculated such that a conditional power of $1-\beta$ is achieved at the observed estimate $\hat{\vartheta}_{1}$ or noncentrality parameter $z_{1}$ of the first stage. There are other strategies for adjusting the information used; see, for example, Section 7.4 in the work of Wassmer and Brannath [15].
@@
We will now describe how two-stage adaptive designs can be used in the two-stage fast-track procedures. For the application for conditional registration, the data from the first stage is used, and for the application for permanent registration, the data from the first and second stages are used. We consider the case where there is no possibility of an early rejection after the first stage, which means we set $\alpha_{1}:=0$. We demand that the conditional error function $A$ fulfills $A(z_{1})\leq 0.5$ for all $z_{1}\in\mathbb{R}$. This ensures that $H_{0}$ can only be rejected after the second stage when $Z_{2}\geq 0$, i.e. a non-negative effect $\hat{\vartheta}_{2}\geq 0$ is observed. After the second stage, the hypothesis $H_{0}$ is rejected iff $Z_{2}\geq\Phi^{-1}(1-A(Z_{1}))$. We will use two different kinds of conditional error functions. The first one is a slight modification of the conditional error function of the inverse normal method with equal weights [31]. The second considered conditional error function is a slight modification of the conditional error function of Fisher’s product test [32] (proposed in the works of Bauer [27] and Bauer and Köhne [28] for use in adaptive designs): We define for $z_{0}\in\mathbb{R}\cup\{-\infty\}$
@@
$A_{F,z_{0}}(z_{1}):=\begin{cases}\frac{c_{F}(z_{0})}{1-\Phi(z_{1})}\wedge 0.5,&\text{if $z_{1}\geq z_{0}$}\\ 0,&\text{otherwise}\end{cases}\penalty 10000\ .$ (12)
The constants $c_{I}(z_{0})$ and $c_{F}(z_{0})$ are chosen such that they are nonnegative and fulfill (with $A:=A_{I,z_{0}}$ or $A:=A_{I,z_{0}}$) condition (10). The chosen $z_{0}$ can be interpreted as a binding futility bound on the z-scale with corresponding early stopping level $\alpha_{0}=1-\Phi(z_{0})$. If there is no possibility for permanent registration after a non-successful conditional registration, then the binding futility boundary $z_{0}=z_{f}$ is a logical choice. This implies that the considered conditional error functions depend on $t_{\xi}(I_{1})$, $\xi$ (and the fixed $\eta_{f}$) because $z_{f}$ depends on these parameters.
@@
$I_{2}^{A}(z_{1}):=I_{2}^{A}(z_{1},I_{2,\min},I_{1}):=\max\left\{I_{2,\min},{I_{1}\left\{\Phi^{-1}(1-\beta)+\Phi^{-1}\big(1-A(z_{1})\big)\right\}^{2}}/{z_{1}^{2}}\right\},\quad z_{1}\geq z_{f},$ (13)
@@
where the minimum information $I_{2,\min}$ remains to be defined. Because $H_{0}$ is tested at the second stage by a z-test at level $A(z_{1})$, we obtain the overall power
@@
$\mathcal{S}_{A}(I_{2,\min},I_{1})$ $:=\mathbb{P}_{\delta}\left(Z_{1}\geq z_{f},\text{reject\penalty 10000\ }H_{0}\right)$
$=\int\limits_{z_{f}}^{\infty}\left\{1-\Phi\left(\Phi^{-1}(1-A(z_{1}))-\eta_{f}{\sqrt{I_{2}^{A}(z_{1})}}/{\sqrt{I_{\delta}}}\right)\right\}\phi\big(z_{1}-\sqrt{I_{1}}\delta\big)dz_{1}\penalty 10000\ .$ (14)
Similar to Subsection 4.1, for achieving an overall power of at least $1-\beta$, the information $I_{1}$ must be strictly greater than $I_{1,\min}$. Furthermore, it can be shown that for $\delta>\delta_{\text{rel}}$ and $\eta_{f}$ and all $I_{1}>I_{1,\min}$ with $\mathcal{S}_{A}(0,I_{1})\leq 1-\beta$ there exists a unique minimum information $I_{2,\min}$ such that $\mathcal{S}_{A}(I_{2,\min},I_{1})=1-\beta$. For $I_{1}$ with $\mathcal{S}_{A}(0,I_{1})\geq 1-\beta$, there is no minimum information needed for the second stage. In the latter case, we set $I_{2,\min}=0$. For proof, we refer to the Appendix. For our example fast-track procedure, the calculation of the second-stage sample size can be done similarly to (9) by replacing $\alpha$ with $A(z_{1})$.
The dashed and solid lines in Figure 3 show the minimum required sample sizes relative to $I_{\delta}$, i.e. $t_{\xi}(I_{2,\min})$, for the two different $\xi$ with the conditional error functions $A_{I,z_{f}}$ and $A_{F,z_{f}}$, respectively. It can be seen that, with regard to the minimum needed information for the second stage, the control of the overall power ($1-\beta=0.8$) at $\delta=\xi\delta_{\text{rel}}$ comes for the adaptive designs with a much smaller requirement for the minimum information than for the non-adaptive design. For the adaptive designs, the minimum second stage information becomes quickly less than the information $I_{\delta}$ required for the fixed-sample size test and is less than 50% of this information if the interim information is $0.55\cdot I_{\delta}$ or more. For the example fast-track procedure, recall that for $\xi=2$ the relative information $t_{\xi}(I_{1})=n_{1}/n_{\delta}=63/105=0.6$ corresponds to the chosen first stage sample size per group $n_{1}=63$. According to Figure 3 (right), the minimum relative information needed for the second stage is now $t_{\xi}(I_{2,\min})=0.33$ and $t_{\xi}(I_{2,\min})=0.29$ for the inverse normal method and Fisher’s product test, respectively. This implies $n_{2,\min}=0.33\cdot n_{\delta}\approx 35$ and $n_{2,\min}=0.29\cdot n_{\delta}\approx 31$, respectively. These sample sizes are much lower than $n_{2,\min}=148$ needed for the non-adaptive design.
4.3 Comparison of different study designs
We now want to compare the efficiency of conducting an adaptive design (with conditional error function $A\leq 0.5$) to a design with two separate stages with regard to the expected and maximum information under the same assumptions as in the previous sections. In particular, this implies that for a given interim information $I_{1}$, the applicant achieves the overall power of $1-\beta$ at the a priori assumed effect $\delta=\xi\cdot\delta_{\text{rel}}$ by choice of the required minimum second stage information $I_{2,\min}$. We will compare the expected second stage information in dependence of $I_{1}>I_{1,\min}$ assuming that the true effect is $\delta$. Like for the first stage, expected and maximum informations will be compared for different $\xi$ and again relative to $I_{\delta}$, i.e. on the $t_{\xi}$-scale for a most general comparison which only depends on $\alpha_{c}$, $\alpha$, $\beta$ and $\xi$.
Figure 4 shows for $\xi=1.75$ and $\xi=2$ the (relative) mean information for the second stage for the two adaptive designs with conditional error functions $A_{I,z_{f}}$ and $A_{F,z_{f}}$ and the non-adaptive design. It can be seen that the adaptive designs have for all $t_{\xi}(I_{1})$ with $t_{\xi}(I_{1,\min})<t_{\xi}(I_{1})\leq t_{\xi}(I_{1,\max})$ a lower (relative) mean information. Furthermore, for the adaptive designs for the most $t_{\xi}(I_{1})$, the relative mean information is (clearly) less than $1$. Thus, the mean information needed for the second stage is less than the information needed for a (single) fixed-size sample test with power $1-\beta$ (direct application for permanent registration). Furthermore, it can be numerically calculated that the mean information over both stages is not much larger than $1$. The corresponding Figures can be found in the Supplement.
Figure 4: Visualization of the mean information for the second stage for $\xi=1.75$ (left) and $\xi=2$ (right) for adaptive designs and a non-adaptive design.
The maximum informations of the considered designs are shown in Figure 5. Figure 4 and Figure 5 have in common that if $t_{\xi}(I_{1})$ comes closer to $t_{\xi}(I_{1,\min})$, the mean and maximum information for the second stage become large and tend to infinity. This is because the minimum information $t_{\xi}(I_{2,\min})$ tends to infinity when $t_{\xi}(I_{1})$ comes closer to the critical value $t_{\xi}(I_{1,\min})$ as described in the previous subsections. As can be seen from Figure 5 (right) for $\xi=2$, the maximum information of all designs has a linear increasing part between $0.45$ and $0.55$. This is because, for $\xi=2$ for all $t_{\xi}(I_{1})\leq 0.55$, the border $z_{f}$ is the same: It holds $z_{f}=1-\Phi^{-1}(1-\alpha_{c})=1.04$ (i.e. the demand $p_{1}\leq\alpha_{c}$ is stricter than having a numeric relevant estimation). This implies that the maximum information is a locally linear function of $t_{\xi}(I_{1})$. For $\xi=1.75$, there is no such interval on which $z_{f}$ is constant. Thus, there is no increasing part in the maximum information in Figure 5 (left). For both adaptive designs for $\xi=1.75$ and $\xi=2$, the relative maximum information remains constant for $t_{\xi}(I_{1})>1.06$ and $t_{\xi}(I_{1})>1.38$, respectively, because the conditional error functions are then constantly equal to $0.5$. Figure 5 shows that the adaptive designs are also convincing with regard to the maximum information when compared with the non-adaptive design. We analyze this result.
Figure 5: Visualization of the maximum information for the second stage for $\xi=1.75$ (left) and $\xi=2$ (right) for adaptive designs and a non-adaptive design.
Recall that the maximum sample size is achieved for $z_{1}=z_{f}$ or equals $I_{2,\min}$. It seems obvious that an adaptive design with conditional error function $A$ has lower maximum information than the non-adaptive design if $A(z_{f})>\alpha$ is fulfilled. Nevertheless, since the maximum information can also be equal to $I_{2,\min}$, a formal proof is needed (see the Appendix). For our considered conditional error functions for $\xi=1.75$ and $\xi=2$ for all $I_{1}>I_{1,\min}$ it holds indeed $A_{I,z_{f}}(z_{f})>\alpha$ and $A_{F,z_{f}}(z_{f})>\alpha$ which gives a formal explanation for the lower maximum informations in Figure 5. It is important to note that the binding futility bound enables the possibility to construct conditional error functions $A$ which fulfill for all $z_{1}\geq z_{f}$ that $A(z_{1})>\alpha$ and thus $A(z_{f})>\alpha$ while having the Type I error controlled at level $\alpha$. In the most extreme case, the conditional error function, which is constant equals ${\alpha}/{\alpha_{f}}>\alpha$, can be chosen with the binding futility bound $Z_{1}<z_{f}$. This design would probably only be accepted if the authorities were informed about it before the start of data collection of the first stage.
Example (continued).
For our example fast-track procedure for the a priori assumed $\xi=2$ and the considered first stage sample size of $n_{1}=63$ (i.e. $t_{\xi}(I_{1})=0.6$) Figure 5 (right) yields $t_{\xi}(I_{2,\max})=4,2.78,3$ for the non-adaptive design, the inverse normal method and Fishers’ product test, respectively. Using the relation $t_{\xi}(I_{2,\max})=n_{2,\max}/n_{\delta}$ and the results from the previous sections we have that the sample size per group for the second stage lies between: $n_{2,\min}=148$ and $n_{2,\max}=420$ (non-adaptive design), $n_{2,\min}=35$ and $n_{2,\max}=292$ (inverse normal method), $n_{2,\min}=31$ and $n_{2,\max}=315$ (Fisher’s product test). Figure 4 (right) yields $t_{\xi}(E(I_{2}))=1.41$ (non-adaptive design) and $t_{\xi}(E(I_{2}))=0.47$ (adaptive designs). This implies a second stage mean sample size per group of $E(n_{2})=1.41\cdot n_{\delta}=149$ (non-adaptive design) and $E(n_{2})=0.47\cdot n_{\delta}\approx 50$ (adaptive designs). The average sample sizes of the second stage are close to the minimum sample sizes, which implies that the distribution of the sample size is left-skewed in the sense that sample sizes close to the maximum ones are rare.
5 Combination of conditional and permanent registration
Recall from Section 3.2 that, under the assumption that the applicants cannot apply for a permanent registration after an unsuccessful attempt to obtain conditional registration, the fast-track procedure does not make sense for all $\xi\leq\xi_{\min}=1.43$. Then, the overall success probability of achieving permanent registration is given by $\mathbb{P}_{\delta}(\text{reject\penalty 10000\ }H_{0})=\mathbb{P}_{\delta}(Z_{1}\geq z_{f},\text{reject\penalty 10000\ }H_{0})$, and we have seen that for $\xi\leq\xi_{\min}=1.43$ there is no first-stage information $I_{1}$ for which an overall success probability of $1-\beta=0.8$ can be achieved, while the requirements for a conditional registration are less strict than those for a direct permanent registration (i.e., $I_{1,\min}>I_{1,\max}$). Motivated by this, we relax the assumption that a conditional registration is a requirement for permanent registration. We assume that, based on the pilot data, we can decide to either apply for conditional registration or abstain from this application and continue directly with the second stage for permanent registration. We will work again with $\alpha_{c}=0.15$. This Section presents possible designs and design inspirations for all a priori effect assumptions $\xi\geq 1$, in particular for an a priori assumed (or true) effect of $\delta/\delta_{\text{rel}}\leq\xi_{\min}$.
5.1 Suggestion for a combination strategy
A reasonable strategy for this is to apply for conditional registration whenever $Z_{1}\geq z_{f}$ (since the requirements are met), and otherwise to abstain from this application and continue the two-step approach with the second stage for permanent registration. This process is illustrated in Figure 6. In order to meet the overall type I error rate $\alpha$, we utilize adaptive design methodology and use for both cases an overall conditional error function $A$ satisfying (10). We will discuss possible choices for $A$ in the following.